研究集会のお知らせ

Workshop on Asymptotic Behavior of Solutions to Partial Differential Equations

日時：１月２５日（木曜日）１０：００−１７：２０：１月２６日（金曜日）１０：００−１７：２０

場所 : 東京理科大学若宮校舎 W-12 教室

平成１３年１月２５日（木曜日）

10:00-10:50 J. M. Delort (Université Paris-Nord)

Quasi-linear Klein-Gordon equations in one space dimension I: global existence and asymptotics

11:00-11:50 A. Hoshiga (Shizuoka University)

The life span of smooth solutions to the systems of quasilinear wave equations with different propagation speeds

昼食

13:30-14:20 H. Hirata (Sophia University)

The Cauchy problem of a semilinear integro-differential equation interporating heat and wave equations

14:30-15:20 M. Nakamura (Tohoku University)

Small data scattering for nonlinear Klein-Gordon equations in the Sobolev spaces

15:30-16:20 S. Tonegawa (Tohoku University)

Global existence for a class of cubic nonlinear Schrödinger equations in one space dimension

16:30-17:20 Y. Yamazaki (Science University of Tokyo) Large time behavior of small solutions to subcritical derivative nonlinear Schrödinger equations

１月２６日（金曜日）

10:00-10:50 J. M. Delort (Université de Paris-Nord)

Quasi-linear Klein-Gordon equations in one space dimension II: energy estimates and applications

11:00-11:50 M. Kadowaki (Tokyo Metropolitan College of Aeronautical Engineering) Resolvent estimates and scattering states for dissipative systems

昼食

13:30-14:20 T. Morioka (Osaka University)

Elastic wave in Shadow

14:30-15:20 N. Kita (Kyushu University)

Sharp asymptotics for derivative nonlinear Schrödinger equations with gauge invariant cubic nonlinearity

15:30-16:20 T. Yokota (Science University of Tokyo)

Remarks on the complex Ginzburg-Landau and nonlinear Schrödinger equations

16:30-17:20 K. Tsugawa (Tohoku University)

Self-similar solutions of the coupled system of semilinear wave equations with different propagation speeds

本研究会は以下の助成によって行われます。文部省科学研究費基盤 A （１）課題番号 ０９３０４０１２「非線型場の大域的構造の函数解析的研究」及び文部省科学研究費基盤 B （２）課題番号 １２４４００５０「非線形分散型方程式の解の性質についての研究」